# Qingdao 266237, China

### v1

@@ 8.3 Experimental Results
LAN 3 LG23 0.030 0.038 0.056 0.106 0.286 1.024 3.899 $\times$ 0.631 0.660 0.822 1.188 4.709 22.62 95.98 $\times$
Our SK 0.004 0.005 0.006 0.012 0.033 0.138 0.749 3.641 0.232 0.254 0.282 0.344 0.824 3.042 14.32 119.2
Our PK 0.061 0.134 0.406 1.466 5.452 22.29 87.23 357.3 0.086 0.163 0.440 1.543 5.753 23.55 92.74 389.6
Our PK 0.083 0.200 0.562 2.039 8.030 31.53 127.4 514.3 0.108 0.229 0.596 2.116 8.331 32.79 133.0 546.6
4 LG23 0.050 0.055 0.063 0.117 0.336 1.274 5.572 $\times$ 0.948 0.970 1.098 1.744 6.939 32.12 135.4 $\times$
Our SK 0.006 0.007 0.008 0.017 0.049 0.202 1.043 4.905 0.442 0.458 0.489 0.665 1.700 7.167 36.16 344.71
Our PK 0.125 0.243 0.734 2.642 10.35 40.83 166.7 -- 0.161 0.280 0.784 2.760 11.00 43.51 178.3 --
Our PK 0.160 0.346 1.033 3.936 15.50 61.19 246.3 -- 0.196 0.383 1.083 4.054 16.16 63.87 257.9 --
5 LG23 0.064 0.066 0.077 0.133 0.383 1.509 7.613 -- 1.325 1.355 1.419 2.680 10.82 48.95 213.6 --
Our SK 0.009 0.010 0.013 0.021 0.061 0.290 1.478 -- 0.677 0.708 0.760 1.114 3.260 13.10 62.35 --
Our PK 0.170 0.401 1.229 4.191 16.41 64.84 262.5 -- 0.211 0.443 1.273 4.348 17.13 67.89 276.3 --
Our PK 0.192 0.509 1.611 6.226 24.62 98.04 400.0 -- 0.233 0.551 1.655 6.383 25.33 101.1 413.8 --
7 LG23 0.099 0.111 0.118 0.233 0.676 3.684 14.45 -- 2.356 2.515 2.752 5.247 23.77 103.6 452.1 --
Our SK 0.014 0.017 0.021 0.040 0.109 0.545 2.701 -- 1.312 1.357 1.510 2.336 8.368 37.17 146.0 --
Our PK 0.273 0.649 2.225 8.325 31.62 125.7 521.2 -- 0.338 0.715 2.314 8.695 33.02 131.7 552.7 --
Our PK 0.339 1.031 3.230 12.25 48.71 195.2 798.1 -- 0.404 1.097 3.319 12.62 50.11 201.3 829.6 --
9 LG23 0.136 0.142 0.18 0.49 1.584 6.121 -- -- 4.169 4.251 4.716 9.418 44.19 191.6 -- --
Our SK 0.025 0.029 0.034 0.058 0.186 0.972 -- -- 2.183 2.272 2.557 5.085 18.93 80.01 -- --
Our PK 0.429 1.064 3.553 13.23 51.74 206.8 -- -- 0.495 1.133 3.660 13.79 54.37 217.9 -- --
Our PK 0.567 1.558 5.346 22.65 80.03 322.5 -- -- 0.633 1.627 5.453 23.21 82.66 333.6 -- --
10 LG23 0.161 0.167 0.215 0.531 2.123 8.301 -- -- 4.943 5.178 5.604 12.82 59.17 252.5 -- --
Our SK 0.03 0.033 0.042 0.073 0.234 1.219 -- -- 2.745 2.883 3.290 6.733 25.12 110.7 -- --
Our PK 0.498 1.338 4.348 16.35 63.54 260.6 -- -- 0.581 1.435 4.492 17.03 66.77 275.4 -- --
Our PK 0.677 1.913 6.552 25.11 98.94 404.3 -- -- 0.760 2.010 6.696 25.79 102.2 419.1 -- --
WAN 3 LG23 4.510 4.516 4.521 4.979 6.295 11.87 17.98 $\times$ 11.87 13.04 15.00 18.33 28.49 85.85 247.6 $\times$
Our SK 2.413 2.570 2.580 3.402 4.129 5.241 10.91 31.83 8.591 9.731 12.81 16.63 25.56 56.39 192.6 826.3
Our PK 4.559 4.952 5.787 7.481 13.91 37.62 131.90 505.93 5.300 5.698 6.694 8.420 15.22 39.86 138.5 538.4
Our PK 4.571 4.662 5.643 7.796 14.82 42.08 150.5 593.6 5.312 5.408 6.550 8.735 16.13 44.33 157.1 626.1
4 LG23 5.789 5.876 6.293 6.956 7.587 13.58 23.32 $\times$ 17.10 19.64 26.64 32.29 49.43 139.9 490.9 $\times$
Our SK 3.376 3.536 3.553 4.408 5.919 8.448 20.65 64.01 11.81 14.41 20.42 27.29 48.65 132.1 499.2 2283
Our PK 6.231 7.181 8.962 11.88 23.57 69.76 251.4 -- 7.374 8.327 10.27 13.22 25.31 73.35 264.1 --
Our PK 6.264 6.436 8.256 12.14 25.92 81.50 294.2 -- 7.407 7.582 9.565 13.48 27.66 85.08 306.9 --
5 LG23 7.556 7.794 7.902 9.102 10.63 15.18 34.82 -- 23.26 26.46 35.49 47.53 83.94 250.6 969.6 --
Our SK 4.501 4.583 5.217 5.955 7.936 12.21 32.14 -- 18.11 21.85 30.93 45.68 91.47 278.9 1112 --
Our PK 8.050 9.745 11.31 16.08 35.07 108.2 388.9 -- 9.593 11.31 13.03 17.84 37.24 112.5 404.4 --
Our PK 8.072 8.869 10.91 10.84 42.83 129.1 468.9 -- 9.615 10.43 12.63 18.85 45.00 133.3 484.4 --
7 LG23 9.722 10.44 11.25 11.67 12.71 22.73 68.51 -- 34.35 45.68 65.99 95.61 203.1 707.2 2881 --
Our SK 7.241 7.414 8.121 9.808 14.52 22.93 69.19 -- 35.65 46.73 62.48 99.11 234.0 817.7 3399 --
Our PK 11.44 14.77 17.55 25.72 61.13 201.6 745.7 -- 13.79 17.14 20.10 28.30 64.28 209.3 779.2 --
Our PK 11.54 14.11 15.89 29.37 70.84 249.3 921.8 -- 13.89 16.48 18.43 31.96 73.98 257.0 955.3 --
9 LG23 11.07 11.41 13.94 14.91 15.86 34.82 -- -- 55.36 74.52 103.4 166.2 413.7 1531 -- --
Our SK 10.62 10.81 11.75 13.88 21.07 39.73 -- -- 57.91 88.27 113.7 189.5 505.3 1881 -- --
Our PK 16.30 19.45 24.87 37.99 93.69 327.8 -- -- 19.45 22.64 28.21 41.33 97.85 341.3 -- --
Our PK 15.65 19.58 22.33 42.38 109.7 396.3 -- -- 18.80 22.76 25.68 45.73 113.8 409.9 -- --
10 LG23 12.81 12.99 16.29 16.83 17.50 45.20 -- -- 66.79 91.10 137.2 126.4 579.0 2205 -- --
Our SK 12.56 12.76 13.60 16.27 25.20 49.94 -- -- 70.02 113.8 140.6 250.8 693.0 2644 -- --
Our PK 19.48 22.71 28.51 43.45 112.5 392.0 -- -- 23.06 26.32 32.28 47.28 117.2 408.5 -- --
Our PK 18.28 21.67 26.72 50.03 134.4 487.3 -- -- 21.87 25.28 30.49 53.86 139.1 503.8 -- --
[…]
m Protocol Set Size n
Online communication (MB) Total communication (MB)
3 LG23 0.157 0.284 0.962 3.662 14.43 57.58 229.8 $\times$ 6.311 7.904 13.37 32.58 114.6 474.3 2054 $\times$
Our SK 0.031 0.111 0.426 1.687 6.778 27.83 112.5 455.0 2.541 3.582 8.529 30.90 132.7 588.7 2615 11533
Our PK 1.854 2.147 3.293 7.984 26.69 102.4 406.8 1634 2.123 2.489 3.711 8.474 27.30 103.0 408.8 1640
Our PK 1.815 1.983 2.655 5.418 16.36 60.78 239.3 959.5 2.084 2.325 3.073 5.908 16.92 61.41 241.3 966.1
4 LG23 0.242 0.449 1.536 5.868 23.15 92.38 368.6 $\times$ 9.591 12.44 23.89 67.25 253.7 1074 4677 $\times$
Our SK 0.060 0.218 0.848 3.376 13.57 55.44 223.9 904.4 3.943 6.399 17.01 67.12 294.7 1315 5851 25819
Our PK 2.781 3.212 4.939 11.98 40.04 153.6 610.2 -- 3.185 3.725 5.566 12.71 40.88 154.5 613.2 --
Our PK 2.723 2.975 3.983 8.130 24.54 91.20 359.0 -- 3.127 3.488 4.610 8.127 25.38 92.12 362.0 --
5 LG23 0.330 0.630 2.173 8.325 32.86 131.2 523.5 -- 13.12 17.75 36.48 110.7 435.2 1869 8172 --
Our SK 0.097 0.361 1.411 5.628 22.63 92.18 372.0 -- 5.511 10.22 30.59 125.3 551.2 2455 10875 --
Our PK 3.708 4.283 6.585 15.97 53.38 204.8 813.6 -- 4.247 4.966 7.421 16.95 54.51 206.0 817.5 --
Our PK 3.630 3.967 5.311 10.84 32.72 121.6 478.7 -- 4.169 4.650 6.147 11.82 33.84 122.8 482.6 --
7 LG23 0.519 1.038 3.635 13.99 55.29 220.8 881.3 -- 19.92 29.90 71.59 243.2 997.9 4327 18930 --
Our SK 0.198 0.750 2.956 11.82 47.55 193.1 778.6 -- 8.853 18.92 65.23 277.2 1233 5505 24402 --
Our PK 5.563 6.425 9.878 23.95 80.07 307.1 1220 -- 6.371 7.450 11.13 25.42 81.76 309.0 1226 --
Our PK 5.445 5.950 7.966 16.25 49.07 182.3 718.0 -- 6.253 6.975 9.220 17.72 50.76 184.2 723.9 --
9 LG23 0.723 1.509 5.347 20.65 81.72 326.3 -- -- 28.15 44.74 115.0 415.3 1742 7609 -- --
Our SK 0.333 1.278 5.060 20.27 81.54 330.5 -- -- 13.90 33.27 120.9 520.0 2310 10268 -- --
Our PK 7.417 8.566 13.17 31.94 106.8 409.5 -- -- 8.495 9.933 14.84 33.90 109.0 412.0 -- --
Our PK 7.26 7.933 10.62 21.67 65.43 243.1 -- -- 8.338 9.300 12.29 23.63 67.68 245.6 -- --
10 LG23 0.831 1.768 6.296 24.36 96.43 385.1 -- -- 32.32 54.34 148.7 549.4 2314 10081 -- --
Our SK 0.413 1.594 6.322 25.34 101.9 412.9 -- -- 16.12 40.87 151.2 656.4 2921 12989 -- --
Our PK 8.344 9.637 14.82 35.93 120.1 460.7 -- -- 9.556 11.18 16.70 38.13 122.6 463.5 -- --
Our PK 8.168 8.925 11.95 24.38 73.61 273.5 -- -- 9.380 10.46 13.83 26.59 76.14 276.3 -- --
[…]
As we can see in the tables, our protocols outperform LG protocol in almost all the case studies. Especially, our SK-MPSU protocol performs best in terms of online running time in LAN setting, online communication costs for a relatively small number of parties, and total running time for medium and large sets in LAN setting among the three protocols. We observe a 4.99.3×4.9-9.3$\times$4.9 - 9.3 × improvement in online running time and a 1.78.5×1.7-8.5$\times$1.7 - 8.5 × improvement in total running time in the LAN setting. It is worth mentioning that our SK-MPSU takes only $3.6$ seconds for 3 parties with sets of $2^{20}$ items each, and $4.9$ seconds for 4 parties with sets of $2^{20}$ items each in online phase, while the implementation of LG protocol fails to run correctly in these cases.
On the other hand, our PK-MPSU protocol performs best in terms of total running time in WAN setting and total communication costs among the three protocols. It shows a 3.036.5×3.0-36.5$\times$3.0 - 36.5 × improvement in overall communication, and a 1.85.4×1.8-5.4$\times$1.8 - 5.4 × improvement in overall running time in the WAN setting, compared to LG protocol.
For the cases where both the number of parties and the set size are relatively large and the network environment is bandwidth-constrained, LG protocol marginally surpasses ours in terms of online running time, but its speedup is only up to 1.4×1.4$\times$1.4 ×. This result aligns with our theoretical analysis in Section 7.2. Nevertheless, when each party’s set size increases to sufficiently large ($2^{20}$), their implementation gives wrong results.

### v4

@@ 7.3 Experimental Results
    3 LG 0.050 0.058 0.069 0.143 0.425 1.582 6.219 25.16 0.982 1.016 1.098 1.767 5.650 32.42 153.8 610.8
    Our SK 0.005 0.007 0.009 0.017 0.050 0.213 1.005 4.352 0.361 0.386 0.395 0.489 1.170 4.534 19.77 96.23
    3 Our PK 0.092 0.237 0.822 3.176 12.69 51.17 205.7 829.1 0.117 0.273 0.868 3.270 13.08 53.05 215.0 872.6
    4 LG 0.069 0.076 0.093 0.162 0.478 1.637 6.375 25.79 1.292 1.358 1.514 2.432 8.716 45.37 197.2 762.9
    Our SK 0.008 0.010 0.013 0.023 0.071 0.286 1.393 5.645 0.639 0.665 0.696 0.917 2.597 10.94 50.25 237.8
    4 Our PK 0.159 0.465 1.570 6.136 24.60 44.91 398.4 1604 0.196 0.505 1.643 6.299 25.45 48.95 417.5 1694
    5 LG 0.107 0.109 0.126 0.190 0.541 1.998 7.588 31.33 1.939 1.993 2.202 3.408 12.54 65.04 289.9 1152
    Our SK 0.012 0.013 0.017 0.030 0.087 0.368 1.714 7.003 0.914 0.964 1.023 1.558 4.551 18.71 85.79 373.4
    5 Our PK 0.232 0.693 2.487 9.858 39.33 158.1 637.5 2816 0.292 0.796 2.547 10.09 40.25 162.2 655.0 2979
    7 LG 0.154 0.165 0.174 0.281 0.795 2.894 10.92 -- 3.484 3.583 3.921 5.84 23.42 111.2 484.5 --
    Our SK 0.019 0.021 0.028 0.048 0.156 0.607 2.817 -- 2.051 2.079 2.214 3.470 12.92 56.50 245.6 --
    7 Our PK 0.463 1.365 5.030 19.43 77.22 310.0 1247 -- 0.550 1.469 5.158 19.73 78.59 316.3 1276 --
    9 LG 0.222 0.226 0.234 0.417 1.216 4.449 17.29 -- 5.501 5.612 6.249 10.27 41.69 182.7 793.3 --
    Our SK 0.027 0.031 0.039 0.075 0.230 0.970 4.293 -- 3.363 3.402 3.895 7.161 30.09 126.3 678.8 --
    9 Our PK 0.764 2.271 8.074 31.64 126.7 507.0 2039 -- 0.880 2.375 8.238 32.19 128.7 515.5 2080 --
    10 LG 0.228 0.243 0.276 0.514 1.479 5.467 -- -- 6.736 6.846 8.335 13.56 55.58 238.3 -- --
    Our SK 0.031 0.036 0.043 0.088 0.286 1.183 -- -- 3.965 4.232 4.821 9.334 41.97 175.9 -- --
    Time. (s) LAN 10 Our PK 0.865 2.800 9.944 39.09 155.9 622.7 2503 -- 0.970 2.912 10.11 39.58 158.4 634.6 2556 --
    3 LG 4.502 4.505 4.522 4.914 6.272 8.744 17.78 54.13 12.46 13.59 15.76 19.52 30.73 78.74 282.8 1188
    Our SK 2.165 2.166 2.332 3.157 3.734 4.444 9.705 33.10 8.403 9.710 12.62 16.30 25.02 56.88 194.7 801.2
    3 Our PK 4.419 4.555 5.553 7.984 18.21 59.77 226.3 900.3 5.168 5.306 6.472 8.968 19.65 62.73 237.3 946.1
    4 LG 5.696 5.880 6.540 7.094 7.323 11.36 23.13 78.81 17.94 20.99 27.63 32.69 50.59 145.7 554.7 2167
    Our SK 2.967 2.969 3.298 3.976 4.618 6.507 17.10 59.21 11.46 13.93 20.21 26.77 47.49 133.2 520.0 2141
    4 Our PK 5.622 5.899 7.929 12.17 32.40 113.8 440.9 1761 6.773 7.052 9.25 13.53 34.27 118.9 461.3 1857
    5 LG 7.385 7.708 8.621 9.198 9.687 14.62 32.83 126.3 23.64 26.82 37.39 48.48 88.32 263.3 1053 4394
    Our SK 3.768 3.733 4.471 4.800 5.521 8.938 25.95 95.40 17.53 21.10 30.28 44.30 88.15 278.9 1119 4820
    5 Our PK 6.849 7.928 10.50 17.08 49.75 179.8 703.3 2873 8.424 9.483 12.22 18.86 52.07 185.4 724.5 2980
    7 LG 9.312 9.833 10.55 11.35 12.14 21.29 66.93 -- 34.92 45.69 66.26 94.89 203.4 705.8 2898 --
    Our SK 5.373 5.381 6.207 6.644 8.164 17.67 56.894 -- 34.00 44.95 61.34 95.03 228.7 817.4 3506 --
    7 Our PK 9.504 12.32 15.14 29.73 92.66 348.5 1377.4 -- 11.88 14.71 17.72 32.35 95.86 356.5 1409 --
    9 LG 11.41 12.21 13.34 14.41 15.09 33.84 115.5 -- 56.65 75.24 104.1 169.1 406.0 1503 6387 --
    Our SK 6.977 7.068 7.830 8.387 12.24 29.20 105.7 -- 58.81 84.80 107.6 182.0 502.5 1915 8137 --
    9 Our PK 13.67 18.84 22.96 45.54 148.2 570.9 -- -- 16.87 22.04 26.31 48.98 152.3 581.5 2034 --
    10 LG 11.77 12.24 15.80 16.49 17.48 45.20 -- -- 66.19 92.22 125.1 219.8 582.1 2179 -- --
    Our SK 7.780 8.032 8.635 9.203 14.54 38.31 -- -- 71.58 109.2 132.7 242.7 684.0 2687 -- --
    Time. (s) WAN 10 Our PK 16.32 23.05 28.66 59.79 184.8 708.5 2792 -- 19.90 26.66 32.44 63.66 189.3 720.9 2846 --
    3 LG 0.157 0.284 0.962 3.662 14.43 57.58 229.8 917.0 6.311 7.904 13.37 32.58 114.6 474.3 2052 8973
    Our SK 0.032 0.111 0.426 1.690 6.788 27.87 112.7 455.6 2.542 3.582 8.529 30.91 132.7 588.8 2614 11529
    3 Our PK 1.815 1.983 2.655 5.418 16.36 60.78 239.3 959.5 2.084 2.325 3.073 5.908 16.92 61.41 240.0 960.3
    4 LG 0.242 0.449 1.536 5.868 23.15 92.38 368.6 1471 9.591 12.44 23.89 67.25 253.7 1074 4674 20419
    Our SK 0.058 0.204 0.791 3.145 12.81 51.69 208.8 843.7 3.941 6.385 16.96 66.89 293.9 1312 5834 25751
    4 Our PK 2.723 2.975 3.983 8.126 24.54 101.9 359.0 1439 3.127 3.488 4.610 8.861 25.38 102.8 360.1 1440
    5 LG 0.330 0.630 2.173 8.325 32.86 131.2 523.5 2090 12.92 17.75 36.48 110.7 435.2 1868 8228 35737
    Our SK 0.090 0.323 1.257 5.007 20.36 82.11 331.5 1339 5.504 10.18 30.44 124.7 548.9 2445 10832 47635
    5 Our PK 3.630 3.967 5.311 10.84 32.72 121.6 478.7 1919 4.169 4.650 6.147 11.82 33.84 122.8 480.1 1921
    7 LG 0.519 1.038 3.635 13.99 55.29 220.8 881.3 -- 19.92 29.90 41.59 243.3 997.9 4326 18924 --
    Our SK 0.172 0.634 2.489 10.03 40.31 162.5 655.4 -- 8.827 18.80 64.76 275.4 1226 5474 24275 --
    7 Our PK 5.445 5.950 7.966 16.25 49.07 182.3 718.0 -- 6.253 6.98 9.220 17.72 50.76 184.2 720.1 --
    9 LG 0.723 1.509 5.347 20.65 81.72 326.3 1303 -- 28.15 44.74 115.0 415.33 1742 7609 33360 --
    Our SK 0.279 1.046 4.122 16.60 66.76 269.0 1084 -- 13.99 33.04 119.9 516.3 2295 10207 45079 --
    9 Our PK 7.260 7.933 10.62 21.67 65.43 243.1 957.3 -- 8.338 9.300 12.29 23.63 67.68 245.6 960.1 --
    10 LG 0.831 1.768 6.296 24.36 96.43 385.1 -- -- 32.32 54.34 148.7 549.4 2314 10081 -- --
    Our SK 0.341 1.288 5.086 20.48 82.39 332.0 -- -- 16.21 40.56 149.9 651.6 2901 12908 -- --
    Comm. (MB) 10 Our PK 8.168 8.925 11.95 24.38 73.61 273.5 1077 -- 9.380 10.46 13.83 26.59 76.14 276.3 1080 --
    Table 2: Online and total running time / communication costs of LG and our protocols in LAN and WAN settings. $m$ is the number of parties. Our SK /PK denotes our SK-MPSU / PK-MPSU protocol. Cells with * denotes trials that the original LG code will give wrong results. Cells with -- denotes trials that ran out of memory. The best protocol within a setting is marked in blue.
    • Acknowledgements. We would like to thank Jiahui Gao for the clarification of their work.
  • References
    • [BA12] Marina Blanton and Everaldo Aguiar. Private and oblivious set and multiset operations. In 7th ACM Symposium on Information, Compuer and Communications Security, ASIACCS 2012, pages 40–41. ACM, 2012.
    • [BC23] Dung Bui and Geoffroy Couteau. Improved private set intersection for sets with small entries. In Public-Key Cryptography - PKC 2023 - 26th IACR International Conference on Practice and Theory of Public-Key Cryptography, volume 13941 of Lecture Notes in Computer Science, pages 190–220. Springer, 2023.
    • [BCG+19a] Elette Boyle, Geoffroy Couteau, Niv Gilboa, Yuval Ishai, Lisa Kohl, Peter Rindal, and Peter Scholl. Efficient two-round OT extension and silent non-interactive secure computation. In Proceedings of the 2019 ACM SIGSAC Conference on Computer and Communications Security, CCS 2019, pages 291–308. ACM, 2019.
    • [BCG+19b] Elette Boyle, Geoffroy Couteau, Niv Gilboa, Yuval Ishai, Lisa Kohl, and Peter Scholl. Efficient pseudorandom correlation generators: Silent OT extension and more. In Advances in Cryptology - CRYPTO 2019. Springer, 2019.
    • [Bea91] Donald Beaver. Efficient multiparty protocols using circuit randomization. In Advances in Cryptology - CRYPTO ’91, 11th Annual International Cryptology Conference, volume 576 of Lecture Notes in Computer Science, pages 420–432. Springer, 1991.
    • [BKM+20] Prasad Buddhavarapu, Andrew Knox, Payman Mohassel, Shubho Sengupta, Erik Taubeneck, and Vlad Vlaskin. Private matching for compute. 2020. https://eprint.iacr.org/2020/599.
    • [BNOP22] Aner Ben-Efraim, Olga Nissenbaum, Eran Omri, and Anat Paskin-Cherniavsky. Psimple: Practical multiparty maliciously-secure private set intersection. In ASIA CCS ’22, pages 1098–1112. ACM, 2022.
    • [BPSY23] Alexander Bienstock, Sarvar Patel, Joon Young Seo, and Kevin Yeo. Near-Optimal oblivious Key-Value stores for efficient PSI, PSU and Volume-Hiding Multi-Maps. In USENIX Security 2023, pages 301–318, 2023.
    • [BS05] Justin Brickell and Vitaly Shmatikov. Privacy-preserving graph algorithms in the semi-honest model. In Advances in Cryptology - ASIACRYPT 2005, 11th International Conference on the Theory and Application of Cryptology and Information Security, volume 3788 of Lecture Notes in Computer Science, pages 236–252. Springer, 2005.
    • [CDG+21] Nishanth Chandran, Nishka Dasgupta, Divya Gupta, Sai Lakshmi Bhavana Obbattu, Sruthi Sekar, and Akash Shah. Efficient linear multiparty PSI and extensions to circuit/quorum PSI. In CCS ’21, pages 1182–1204. ACM, 2021.
    • [CGP20] Melissa Chase, Esha Ghosh, and Oxana Poburinnaya. Secret-shared shuffle. In Advances in Cryptology - ASIACRYPT 2020, volume 12493 of Lecture Notes in Computer Science, pages 342–372. Springer, 2020.
    • [CGS22] Nishanth Chandran, Divya Gupta, and Akash Shah. Circuit-psi with linear complexity via relaxed batch OPPRF. Proc. Priv. Enhancing Technol., 2022(1):353–372, 2022.
    • [CM20] Melissa Chase and Peihan Miao. Private set intersection in the internet setting from lightweight oblivious PRF. In Advances in Cryptology - CRYPTO 2020, volume 12172 of Lecture Notes in Computer Science, pages 34–63. Springer, 2020.
    • [CO18] Michele Ciampi and Claudio Orlandi. Combining private set-intersection with secure two-party computation. In Security and Cryptography for Networks - 11th International Conference, SCN 2018, volume 11035 of Lecture Notes in Computer Science, pages 464–482. Springer, 2018.
    • [CZZ+24a] Yu Chen, Min Zhang, Cong Zhang, Minglang Dong, and Weiran Liu. Private set operations from multi-query reverse private membership test. In Public-Key Cryptography - PKC 2024. Springer, 2024.
    • [CZZ+24b] Yu Chen, Min Zhang, Cong Zhang, Minglang Dong, and Weiran Liu. Private set operations from multi-query reverse private membership test. In PKC 2024, 2024. https://eprint.iacr.org/2022/652.
    • [EB22] Saba Eskandarian and Dan Boneh. Clarion: Anonymous communication from multiparty shuffling protocols. In 29th Annual Network and Distributed System Security Symposium, NDSS 2022. The Internet Society, 2022.
    • [ElG85] Taher ElGamal. A public key cryptosystem and a signature scheme based on discrete logarithms. IEEE Transactions on Information Theory, 31:469–472, 1985.
    • [FIPR05] Michael J. Freedman, Yuval Ishai, Benny Pinkas, and Omer Reingold. Keyword search and oblivious pseudorandom functions. In Theory of Cryptography, Second Theory of Cryptography Conference, TCC 2005, volume 3378 of Lecture Notes in Computer Science, pages 303–324. Springer, 2005.
    • [Fri07] Keith B. Frikken. Privacy-preserving set union. In Applied Cryptography and Network Security, 5th International Conference, ACNS 2007, volume 4521 of Lecture Notes in Computer Science, pages 237–252. Springer, 2007.
    • [GMR+21] Gayathri Garimella, Payman Mohassel, Mike Rosulek, Saeed Sadeghian, and Jaspal Singh. Private set operations from oblivious switching. In Public-Key Cryptography - PKC 2021, volume 12711 of Lecture Notes in Computer Science, pages 591–617. Springer, 2021.
    • [GMW87] Oded Goldreich, Silvio Micali, and Avi Wigderson. How to play any mental game or A completeness theorem for protocols with honest majority. In Proceedings of the 19th Annual ACM Symposium on Theory of Computing, 1987, pages 218–229. ACM, 1987.
    • [GNT23] Jiahui Gao, Son Nguyen, and Ni Trieu. Toward a practical multi-party private set union. Cryptology ePrint Archive, Paper 2023/1930, 2023. Version: 20240316:210303, https://eprint.iacr.org/archive/2023/1930/20240316:210303.
    • [GPR+21] Gayathri Garimella, Benny Pinkas, Mike Rosulek, Ni Trieu, and Avishay Yanai. Oblivious key-value stores and amplification for private set intersection. In Advances in Cryptology - CRYPTO 2021, volume 12826 of Lecture Notes in Computer Science, pages 395–425. Springer, 2021.
    • [HLS+16] Kyle Hogan, Noah Luther, Nabil Schear, Emily Shen, David Stott, Sophia Yakoubov, and Arkady Yerukhimovich. Secure multiparty computation for cooperative cyber risk assessment. In IEEE Cybersecurity Development, 2016, pages 75–76. IEEE Computer Society, 2016.
    • [JSZ+22] Yanxue Jia, Shi-Feng Sun, Hong-Sheng Zhou, Jiajun Du, and Dawu Gu. Shuffle-based private set union: Faster and more secure. In USENIX 2022, 2022.
    • [JSZG24] Yanxue Jia, Shi-Feng Sun, Hong-Sheng Zhou, and Dawu Gu. Scalable private set union, with stronger security. In 33rd USENIX Security Symposium, USENIX Security 2024. USENIX Association, 2024.
    • [KKRT16] Vladimir Kolesnikov, Ranjit Kumaresan, Mike Rosulek, and Ni Trieu. Efficient batched oblivious PRF with applications to private set intersection. In CCS 2016, pages 818–829. ACM, 2016.
    • [KMP+17] Vladimir Kolesnikov, Naor Matania, Benny Pinkas, Mike Rosulek, and Ni Trieu. Practical multi-party private set intersection from symmetric-key techniques. In Proceedings of the 2017 ACM SIGSAC Conference on Computer and Communications Security, CCS 2017, pages 1257–1272. ACM, 2017.
    • [KRTW19] Vladimir Kolesnikov, Mike Rosulek, Ni Trieu, and Xiao Wang. Scalable private set union from symmetric-key techniques. In Advances in Cryptology - ASIACRYPT 2019, volume 11922 of Lecture Notes in Computer Science, pages 636–666. Springer, 2019.
    • [KS05] Lea Kissner and Dawn Xiaodong Song. Privacy-preserving set operations. In Advances in Cryptology - CRYPTO 2005, volume 3621 of Lecture Notes in Computer Science, pages 241–257. Springer, 2005.
    • [LG23] Xiang Liu and Ying Gao. Scalable multi-party private set union from multi-query secret-shared private membership test. In Advances in Cryptology - ASIACRYPT 2023. Springer, 2023.
    • [LV04] Arjen K. Lenstra and Tim Voss. Information security risk assessment, aggregation, and mitigation. In Information Security and Privacy: 9th Australasian Conference, ACISP 2004., volume 3108 of Lecture Notes in Computer Science, pages 391–401. Springer, 2004.
    • [MS13] Payman Mohassel and Seyed Saeed Sadeghian. How to hide circuits in MPC an efficient framework for private function evaluation. In Advances in Cryptology - EUROCRYPT 2013, volume 7881 of Lecture Notes in Computer Science, pages 557–574. Springer, 2013.
    • [NTY21] Ofri Nevo, Ni Trieu, and Avishay Yanai. Simple, fast malicious multiparty private set intersection. In Proceedings of the 2021 ACM SIGSAC Conference on Computer and Communications Security, CCS 2021, pages 1151–1165. ACM, 2021.
    • [PR04] Rasmus Pagh and Flemming Friche Rodler. Cuckoo hashing. Journal of Algorithms, 51(2):122–144, 2004.
    • [PRTY19] Benny Pinkas, Mike Rosulek, Ni Trieu, and Avishay Yanai. Spot-light: Lightweight private set intersection from sparse OT extension. In Advances in Cryptology - CRYPTO 2019, volume 11694 of Lecture Notes in Computer Science, pages 401–431. Springer, 2019.
    • [PRTY20] Benny Pinkas, Mike Rosulek, Ni Trieu, and Avishay Yanai. PSI from paxos: Fast, malicious private set intersection. In Advances in Cryptology - EUROCRYPT 2020, volume 12106 of Lecture Notes in Computer Science, pages 739–767. Springer, 2020.
    • [PSSZ15] Benny Pinkas, Thomas Schneider, Gil Segev, and Michael Zohner. Phasing: Private set intersection using permutation-based hashing. In 24th USENIX Security Symposium, USENIX Security 15, Washington, D.C., USA, August 12-14, 2015, pages 515–530. USENIX Association, 2015.
    • [PSTY19] Benny Pinkas, Thomas Schneider, Oleksandr Tkachenko, and Avishay Yanai. Efficient circuit-based PSI with linear communication. In Advances in Cryptology - EUROCRYPT 2019, volume 11478 of Lecture Notes in Computer Science, pages 122–153. Springer, 2019.
    • [PSZ14] Benny Pinkas, Thomas Schneider, and Michael Zohner. Faster private set intersection based on OT extension. In Proceedings of the 23rd USENIX Security Symposium, 2014, pages 797–812. USENIX Association, 2014.
    • [Rab05] Michael O. Rabin. How to exchange secrets with oblivious transfer. IACR Cryptol. ePrint Arch., page 187, 2005.
    • [Roy22] Lawrence Roy. Softspokenot: Quieter OT extension from small-field silent VOLE in the minicrypt model. In Advances in Cryptology - CRYPTO 2022. Springer, 2022.
    • [RR22] Srinivasan Raghuraman and Peter Rindal. Blazing fast PSI from improved OKVS and subfield VOLE. In ACM CCS 2022, 2022.
    • [RRT23] Srinivasan Raghuraman, Peter Rindal, and Titouan Tanguy. Expand-convolute codes for pseudorandom correlation generators from LPN. In Advances in Cryptology - CRYPTO 2023, volume 14084 of Lecture Notes in Computer Science, pages 602–632. Springer, 2023.
    • [RS21] Peter Rindal and Phillipp Schoppmann. VOLE-PSI: fast OPRF and circuit-psi from vector-ole. In Advances in Cryptology - EUROCRYPT 2021, volume 12697 of Lecture Notes in Computer Science, pages 901–930. Springer, 2021.
    • [VCE22] Jelle Vos, Mauro Conti, and Zekeriya Erkin. Fast multi-party private set operations in the star topology from secure ands and ors. IACR Cryptol. ePrint Arch., page 721, 2022.
    • [ZCL+23] Cong Zhang, Yu Chen, Weiran Liu, Min Zhang, and Dongdai Lin. Optimal private set union from multi-query reverse private membership test. In USENIX 2023, 2023. https://eprint.iacr.org/2022/358.
    • [ZMS+21] Shengnan Zhao, Ming Ma, Xiangfu Song, Han Jiang, Yunxue Yan, and Qiuliang Xu. Lightweight threshold private set intersection via oblivious transfer. In Wireless Algorithms, Systems, and Applications - 16th International Conference, WASA 2021, volume 12939 of Lecture Notes in Computer Science, pages 108–116. Springer, 2021.